# Cooperative and Asynchronous Transformer-based Mission Planning for Heterogeneous Teams of Mobile Robots

### v1

@@ Abstract
Coordinating heterogeneous teams of mobile robots for tasks such as search and rescue is highly challenging. This is due to the complexities of perception, decision making and planning in such environments, with agents’ non-synchronous operation, constrained communication, and limited computational resources. This paper presents the Cooperative and Asynchronous Transformer-based Mission Planning (CATMiP) framework, which leverages multi-agent reinforcement learning (MARL) to effectively coordinate agents with heterogeneous sensing, motion, and actuation capabilities. The framework introduces a Class-based Macro-Action Decentralized Partially Observable Markov Decision Process (CMD-POMDP) model to handle asynchronous decision-making among different agent classes via macro-actions. It also extends the Multi-Agent Transformer (MAT) architecture to facilitate distributed, ad hoc communication among the agents. CATMiP easily adapts to mission complexities and communication constraints, and scales to varying environment sizes and team compositions. Simulations demonstrate its scalability and ability to achieve cooperative mission objectives with two classes of explorer and rescuer agents, even under severe communication constraints. The code is available at https://github.com/mylad13/CATMiP.
Index Terms:
Reinforcement Learning, Multi-Robot Systems, Motion and Path Planning, Search and Rescue Robots, Task and Motion Planning, Cooperating Robots.
I INTRODUCTION
Multi-robot systems (MRS) are becoming increasingly prevalent in search and rescue operations [1], monitoring and inspection [2], and industrial automation [3], where coordinated efforts improve efficiency and adaptability in complex tasks. Particularly, heterogeneous systems composed of agents with complimentary capabilities can be more effective than homogeneous teams in carrying out complex missions [4, 1, 5]. However, leveraging the full potential of heterogeneous MRS requires scalable and robust distributed control strategies that can enable seamless coordination among robots with varying roles and abilities, while maintaining resilience to failures and minimizing communication overhead.
Multi-agent reinforcement learning (MARL) allows robots to autonomously learn coordination strategies through interaction with their environment [6]. However, MARL is known to suffer from expensive computations and the reality gap when transferring from simulation to real-world deployment [7, 8]. Additionally, MARL typically assumes synchronous agent actions, whereas in many real-world scenarios robotic agents operate asynchronously [9].
This paper proposes Cooperative and Asynchronous Transformer-based Mission Planning (CATMiP), an MARL framework designed to train heterogeneous teams of mobile robots to effectively cooperate with sporadic communication and asynchronous decision-making. CATMiP introduces a Class-based Macro-Action Decentralized Partially Observable Markov Decision Process (CMD-POMDP) to model asynchronous operation of heterogeneous teams through macro-actions. It extends the Multi-Agent Transformer (MAT) architecture [10] to support asynchronous and distributed decision-making via ad hoc and sporadic communication. CATMiP is scalable and robust with respect to the map size, team size, and team composition throughout the mission. It also uses domain and action delay randomization to help bridge the reality gap.
The rest of the paper is organized as follows. Section II reviews prior relevant work. Section III states and formulates the problem of mission planning for heterogeneous robotic teams along with the proposition of CMD-POMDP. Section IV describes the CATMiP framework in detail. Section V describes the simulation environment set-up, including the observation and action spaces and the reward structure. The results of the simulations are presented and discussed in Section VI. The paper is concluded in Section VII, with a brief discussion of potential directions for future work.
II Background Literature
@@
Deep reinforcement learning (DRL) has been increasingly used in mobile robotics for exploration and navigation, handling complex tasks in both single-agent and multi-agent settings [11, 6, 12]. Such control strategies are typically divided into end-to-end and two-stage approaches. The end-to-end methods derive control actions directly from sensor data, whereas the two-stage approaches first select target locations using RL and then employ a separate method for control actions, improving sample efficiency and generalization. In single-agent scenarios, DRL-based two-stage planning combines high-level goal selection with classical path-planning algorithms [13, 14, 15, 16, 17]. Deep multi-agent reinforcement learning (DMARL) methods have been developed to enhance exploration efficiency and handle communication dropouts in multi-robot settings [18, 19, 20]. DMARL solutions for coordinating heterogeneous agents have also been explored in recent works, such as [21, 22, 23, 24].
Asynchronous reinforcement learning approaches such as Asynchronous Coordination Explorer and Qrainbow have been used to make real-time decisions in the presence of communication delay [9, 25]. They utilize decentralized execution and attention-based relations encoding to reduce communication traffic and improve generalization in dynamic environments.
There are several challenges in the way of applying deep multi-agent reinforcement learning for complex cooperative tasks including large computations, nonstationarity, partial observability, and credit assignment [7]. The Centralized Training and Decentralized Execution (CTDE) scheme [26] attempts to overcome some of these hurdles. CTDE balances computational cost and coordination by allowing agents to share information during training, while requiring them to act on their local observations during execution. Some methods share the policy network parameters among the agents to improve sample efficiency [27, 28]. However, this leads to homogeneous behaviour and sub-optimal performance in complex tasks that require behavioral diversity among the agents. Additionally, intrinsic rewards [29] and value function factorization [30] have proven effective in improving sample efficiency and solving the credit assignment problem.
The Multi-Agent Advantage Decomposition Theorem [31] shows that MARL problems can be decomposed into multiple sequential RL problems. This theorem suggests that agents can achieve a positive joint advantage by sequentially selecting their local actions, rather than searching through the entire joint action space simultaneously. By framing MARL as a sequence modeling problem, its complexity is reduced from multiplicative to additive [10], enabling faster training and more effective handling of heterogeneous teams.
Transformers [32], known for their success in natural language processing and other domains [33], are well-suited for sequence modeling due to their ability to capture relationships within sequences using scaled dot-product attention. This mechanism calculates attention weights by taking the dot product between a matrix of queries ($Q\in\mathbb{R}^{n\times d_{k}}$) and keys ($K\in\mathbb{R}^{n\times d_{k}}$), scaling the result, and normalizing it by a softmax function,
$Attention(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (1)
where $n$ is the length of the input sequence and $d_{k}$ is the dimension of the queries and keys.
The weighting mechanism allows the model to focus on the most relevant parts of the input sequence by applying the calculated attention weights to the values ($v_{i}\in\mathbb{R}^{d_{v}}$). Transformers and the attention mechanism have been recently applied in RL [34]. Particularly, the Multi-Agent Transformer (MAT) architecture [10] has achieved state-of-the-art performance in cooperative MARL tasks, by enabling the agents to learn unique behaviors and focus on critical parts of each other’s observations, enhancing their coordination. The present work extends MAT to accommodate macro-actions and asynchronous updates, leveraging the transformer’s ability to handle variable sequence lengths.
III Problem Statement and Formulation
Figure 1: (a) The centralized training workflow of the CATMiP framework in a simulated mission episode. (b) Implementation of the CATMiP framework on a single robot during the execution phase, where robots share mapping information and their embedded macro-observations to generate new actions in a distributed way.
@@ III Problem Statement and Formulation
This paper is concerned with the design of distributed controllers for a heterogeneous team of mobile robots executing a cooperative mission in an unknown environment. In particular, it focuses on an indoor search and target acquisition scenario with two explorer and rescuer agent types. The mission requires a rescuer agent to reach a target, whose location is initially unknown, as quickly as possible. The control policies must leverage the diverse capabilities of the team, encouraging specialized behaviors for each agent class. For example, the explorer robots are faster and better suited to rapidly mapping the environment and finding the target, whereas a rescuer is slower but is capable of engaging with the target.
The robots use a combination of global and local planning for navigation; the global planner selects a goal within an area centered on the robot and computes the shortest path to it, which serves as the robot’s macro-action (MA), and the local planner generates motion commands to handle obstacle avoidance and smooth movement along the path.
@@
This work introduces a novel formulation for the two-level asynchronous control problem in heterogeneous teams of agents carrying out a cooperative mission. This is accomplished by defining a Class-based Macro Action Decentralized Partially Observable Markov Decision Process (CMD-POMDP). It is an extension of the MacDec-POMDP model [35] that supports varying properties across different agent classes. Assuming a team of $N$ agents with $C\leq N$ different classes, the problem is formalized as the tuple $\langle I,C,S,\{M^{c}\},\{A^{c}\},T,\{R^{c}\},\allowbreak\{\zeta^{c}\},\{Z^{c}\},\{\Omega^{c}\},\{O^{c}\},\gamma,h\rangle$, where
@@
  • $M^{c}$ is a finite set of MAs for agents of class $c$. The set of joint MAs is then $M=\times_{i}M^{C(i)}$;
@@
  • $\zeta^{c}$ is a finite set of macro observations (MOs) for agents of class $c$. The set of MOs is then $\zeta=\times_{i}\zeta^{C(i)}$;
@@
The solution to the CMD-POMDP is the joint high-level policy $\bar{\mu}=(\mu^{1},\dots,\mu^{N})$ that maximizes the expected cumulative discounted return from all agents,
@@
It should be noted that a generic parameter $p$ is denoted as $p^{i}$ when referring to the $i$-th agent, and as $\bar{p}=(p^{1},p^{2},\dots,p^{N})$ when referring to the team of $N$ agents.
IV Proposed Framework
The proposed CATMiP framework for asynchronous and distributed mission planning in cooperative heterogeneous robotic teams follows the CTDE principle. It consists of several key modules: the Asynchronous Multi-Agent Transformer (AMAT), Collaborative Simultaneous Localization and Mapping (CSLAM), the High-Level Processor, and the Low-Level Navigation Controller. These modules are implemented and utilized differently during the training and execution phases, as described later in this section.
Fig. 1 (a) depicts the CATMiP architecture and data flow at time step $t$ of a simulated training episode with full communication, where the active subset $i_{1:g}$ of the agents are obtaining new MAs. During mission execution, each agent generates MAs individually or together with other active agents in its communication neighborhood, as shown in Fig. 1 (b).
@@ IV Proposed Framework
At each time step of the mission, agents are considered active, inactive, or on standby. Inactive agents continue executing primitive actions generated by their current MA but enter standby mode when their MA termination condition is met. Agents on standby remain at their current location and delay selecting their next MA by a random number of time steps. This delay not only allows them to find other agents on standby within their communication neighborhood so they can activate simultaneously and obtain their new MAs with one pass through the network, but it also helps simulate the variable time it might take to execute an MA in real-world scenarios. If an agent’s maximum standby time is reached and no other on-standby agents are found in the neighborhood, it activates alone. Agents that update their MAs become inactive again.
IV-A Centralized Training
@@
The pseudocode of the centralized training process is presented in Algorithm 1. At every time step $t$ of a training episode, the agents perform the primitive actions generated by the navigation controller module based on each agent’s current MA and collect the state-action-dependent reward. At the $\tau$th instance of there being a subset $i_{1:g(\tau)}$ of agents ready to update their MAs (flagged as active agents), an ordered sequence of macro-observations (MOs) $\bar{z}_{\tau}=\{z^{i_{1}},\dots,z^{i_{g(\tau)}},z^{i_{g(\tau)+1}},\dots,z^{i_{N}}\}$ are formed and fed as the input to the AMAT module. Subsequently, a sequence of latent representation of MOs of active agents, $\hat{z}^{i_{1:g(\tau)}}_{\tau}$, are calculated and used to obtain new MAs $m^{i_{1:g(\tau)}}_{\tau}$ and value function estimates $V(\hat{z}^{i_{1:g(\tau)}}_{\tau})$. The rewards accumulated by the active agents since their previous activation $R^{i_{1:g(\tau)}}_{acc}$, along with $\bar{z}_{\tau}$, $m^{i_{1:g(\tau)}}_{\tau}$, and $V(\hat{z}^{i_{1:g(\tau)}}_{\tau})$ are stored in a buffer to be used for updating the network parameters at the end of each episode. The fully trained network is then sent to each agent to be used for inference during the distributed execution phase.
@@
IV-B Distributed Execution
During the execution phase, the robots establish time-varying communication neighborhoods through a dynamic mobile ad hoc network. This setup is modeled by a communication graph $\mathcal{G}=(\mathcal{I},\mathcal{E})$, where the vertices $\mathcal{I}$ represent the set of agents and the edges $\mathcal{E}$ represent the communication links between them. The probability of a communication link $E_{ij}$ existing between agents $i$ and $j$ is modeled as $Pr(E_{ij})=e^{-d_{ij}^{2}/\sigma^{2}}$, where $d_{ij}$ denotes the distance between agents, and $\sigma$ controls the rate of probability decay. This model is adapted from [36]. An agent’s neighborhood, $\mathcal{N}(i)$, consists of all agents (including itself) that can exchange information either directly or indirectly with agent $i$.
Agents within a neighborhood that have been activated at the same time step elect a broker to handle the generation and transmission of new MAs. The broker forms an ordered sequence of MO embeddings $(z_{e}^{i_{1}},\dots,z_{e}^{i_{m}})$, starting with its own MO embedding $z_{e}^{i_{1}}$ and followed by $z_{e}^{i_{2:g}}$ received from active and $z_{e}^{i_{{g+1}:m}}$ received from inactive agents in the neighborhood. This sequence is fed as input to the broker’s local AMAT network, generating a sequence of new MAs $(m^{i_{1}},\dots,m^{i_{g}})$ to be sent to their respective agent.
The following subsections detail each module’s role in the training and execution phases.
IV-C Asynchronous Multi-Agent Transformer
The AMAT module, illustrated in Fig. 2, consists of four components: the macro-observations embedder, the encoder, the macro-actions embedder, and the decoder. In summary, AMAT processes a sequence of macro-observations from agents $i_{1:m}$, accompanied by a vector of active flags indicating the subset $i_{1:g}$ ($g\leq m$) of these agents requiring new macro-actions. It then outputs the corresponding sequence of MAs. The input and output sequence lengths are flexible, allowing AMAT to model policies for a variable number of agents within a single network, and to support asynchronous and distributed inference.
During centralized training, $i_{1:m}$ represents the complete agent set $I=\{1,\dots,N\}$, while in the distributed execution phase, it denotes the set of agents in the same neighborhood $\mathcal{N}(i)=\{i_{1},\dots,i_{m}\}$. In both cases, the sets are ordered to put the active agents $i_{1:g}$ first.

### v2

@@ Abstract
Cooperative mission planning for heterogeneous teams of mobile robots presents a unique set of challenges, particularly when operating under communication constraints and limited computational resources. To address these challenges, we propose the Cooperative and Asynchronous Transformer-based Mission Planning (CATMiP) framework, which leverages multi-agent reinforcement learning (MARL) to coordinate distributed decision making among agents with diverse sensing, motion, and actuation capabilities, operating under sporadic ad hoc communication. A Class-based Macro-Action Decentralized Partially Observable Markov Decision Process (CMacDec-POMDP) is also formulated to effectively model asynchronous decision-making for heterogeneous teams of agents. The framework utilizes an asynchronous centralized training and distributed execution scheme that is developed based on the Multi-Agent Transformer (MAT) architecture. This design allows a single trained model to generalize to larger environments and accommodate varying team sizes and compositions. We evaluate CATMiP in a 2D grid-world simulation environment and compare its performance against planning-based exploration methods. Results demonstrate CATMiP’s superior efficiency, scalability, and robustness to communication dropouts, highlighting its potential for real-world heterogeneous mobile robot systems. The code is available at https://github.com/mylad13/CATMiP.
keywords:
Autonomous Robots , Motion Planning , Robot Learning , Reinforcement Learning.
journal: Robotics and Autonomous Systems
[…]
organization=Department of Electrical and Computer Engineering, McMaster University,addressline=1280 Main St W, city=Hamilton, postcode=L8S 4L8, state=ON, country=Canada
1 INTRODUCTION
Multi-robot systems (MRS) are becoming increasingly prevalent in applications such as search and rescue operations [1], environmental monitoring [2], building and infrastructure inspection, [3], and industrial plant management [4]. The coordinated efforts of robots in these systems improve efficiency and adaptability, especially in complex tasks. Particularly, heterogeneous MRS composed of robots with complimentary capabilities outperform homogeneous teams in missions requiring diverse sensing and actuation capabilities [5, 6, 7].
Coordination in MRS can be centralized or decentralized. Centralized approaches rely on a leader robot or server to issue commands, which can result in high computational and communication loads, vulnerability to single-point failures, and challenges in ensuring consistent communication in real-world scenarios. In contrast, decentralized approaches allow robots to individually make autonomous decisions while implicitly considering the actions of others and changes in the environment. Combining implicit coordination with explicit communication through ad hoc wireless mesh networks enables distributed control strategies that are both scalable and efficient [8, 9].
Deep Multi-Agent Reinforcement Learning (MARL) has emerged as a powerful tool for coordinating MRS in dynamic and uncertain environments [10]. Deep MARL enables robots to learn coordination strategies autonomously, bypassing the need for predefined algorithms and heuristics. However, conventional MARL often assumes synchronous decision-making, where agents take new actions at the same time—a condition that is inefficient and impractical for many real-world scenarios. Asynchronous MARL [11] addresses this limitation by enabling agents to make decisions over temporally extended actions, otherwise known as macro-actions [12].
This paper addresses the distributed coordination of heterogeneous mobile robots navigating unknown environments by proposing the Cooperative and Asynchronous Transformer-based Mission Planning (CATMiP) framework. CATMiP is formulated based on the Class-based Macro-Action Decentralized Partially Observable Markov Decision Process (CMacDec-POMDP) model, a novel extension of the MacDec-POMDP model [12] for decentralized multi-agent planning that considers varying properties across different agent classes. Our case study involves two robot types-explorers and rescuers, where the objective is for a rescuer type robot to reach a target with an initially unknown location as fast as possible.
The robots perform collaborative simultaneous localization and mapping (C-SLAM), merging local occupancy grid maps into a shared global map through intermittent communication [13]. Navigation decisions are made in a distributed manner and following a hierarchical two-level control approach. At the high-level, each robot selects a macro-action, which is a goal point location on the map within a fixed distance from the robot. This macro-action is sampled from a policy generated by the proposed Asynchronous Multi-Agent Transformer (AMAT) network. The inputs to AMAT are the agents’ macro-observations, which are class-specific multi-channeled global and local maps. At the low-level, path planning and motion control modules generate the robot’s immediate action to navigate toward the selected goal.
Section 2 reviews related works and discusses how the identified research gaps are addressed in this paper. Section 3 introduces the CMacDec-POMDP formulation, and formally states the problem. Section 4 details the different components of the CATMiP framework, as well as the asynchronous training and execution phases of AMAT. Section 6 describes the simulation setup, including macro-observations, macro-actions, and reward structures. Simulation results are presented and analyzed in Section 7. Finally, Section 8 concludes the paper and outlines future directions.
2 Related Works
@@
Deep reinforcement learning (DRL) has been increasingly used in mobile robotics for exploration and navigation, handling complex tasks in single-agent and multi-agent settings [14, 10, 15]. Such control strategies are typically divided into end-to-end and two-stage approaches. The end-to-end methods derive control actions directly from sensor data, whereas the two-stage approaches first select target locations using DRL and then employ a separate method for control actions, improving sample efficiency and generalization. Notable recent works have combined high-level DRL-based goal selection with classical path-planning algorithms in single robot scenarios [16, 17, 18].
Cooperative multi-robot mission planning has been studied using various deep MARL approaches. Notable works addressing asynchronous multi-robot exploration with homogeneous robots and macro-actions include [19, 20]. Tan et al. [19] tackle the challenge of communication dropouts in multi-robot exploration by modeling the problem as a MacDec-POMDP and proposing a DRL solution based on the centralized training and decentralized execution (CTDE) paradigm [21]. CTDE strikes a balance between coordination and scalability by enabling agents to learn from shared experiences during training while acting independently based on local observations during execution. Yu et al. [20] extend the multi-agent proximal policy optimization (MAPPO) algorithm [22] to enable asynchronous CTDE. Their approach enhances coordination efficiency through an attention-based relation encoder, which aggregates feature maps from different agents to capture intra-agent interactions. While these methods demonstrate the effectiveness of macro-actions in asynchronous decision-making, all agents follow the same policy and the unique challenges of planning for heterogeneous multi-robot systems are not considered.
To address heterogeneity, Zhang et al. [23] propose an architecture for asynchronous multi-robot decision-making that combines value function decomposition [24], the MacDec-POMDP framework, and the CTDE paradigm. Their approach utilizes features extracted from both global states and local observations during training. However, during execution, each agent generates macro-actions based solely on feature maps derived from its local observations. This design enables diverse behaviors among agents but restricts the trained model to be used by a fixed team size and composition. Moreover, none of the aforementioned methods allow the trained models to generalize to larger environments. This limitation highlights the need for approaches that prioritize scalability and adaptability in multi-robot mission planning.
To enable agent heterogeneity while maintaining the benefits of parameter sharing, agent indication was formalized in [25]. This method appends an agent-specific indicator signal to the observations, allowing a shared policy network to generate agent-specific actions. Terry et al. [25] demonstrated that parameter sharing can be effectively applied to heterogeneous observation and action spaces while still achieving optimal policies. This idea is used in the Multi-Agent Transformer (MAT) architecture [26] as well, where positional encoding that appears in the original transformer [27] are replaced by agent indication.
Wen et al. [26] introduced MAT alongside a novel MARL training paradigm that achieves linear time complexity and guarantees monotonic performance improvement by leveraging the multi-agent advantage decomposition theorem [28]. This theorem suggests that joint positive advantage can be achieved by sequentially selecting local actions rather than searching the entire joint action space simultaneously. Thus, cooperative MARL can be reformulated as a sequence modeling problem, where the objective is to map a sequence of agent observations to a sequence of optimal agent actions.
In MAT, the attention mechanism [27] in the encoder captures the inter-agent relationships within the sequence of observations, and the decoder auto-regressively generates actions by considering the input sequence’s latent representation. The transformer model’s ability to process flexible sequence lengths enables generalization to different team sizes without treating varying agent numbers as separate tasks. This property allows a single trained model to scale to teams with more or less agents than those encountered during training.
Building on MAT, our work introduces the Asynchronous Multi-Agent Transformer (AMAT) network. We develop a new asynchronous centralized training and asynchronous distributed execution scheme tailored for heterogeneous teams. Specialized agent class policies are learned through agent class encodings, which differentiate macro-observations across agent classes and enable the network to generate corresponding macro-actions. This design enhances the model’s generalizability to larger teams with varying compositions of heterogeneous agents. Additionally, we employ an adaptive pooling layer during global feature extraction from macro-observations, allowing the model to efficiently scale to larger environment sizes without compromising performance.
3 Problem Statement and Formulation
@@
This paper addresses the design of distributed controllers for a heterogeneous team of mobile robots performing a cooperative mission in an unknown environment. Specifically, we focus on an indoor search and target acquisition scenario involving two agent classes: explorers and rescuers. The mission objective is for a rescuer agent to reach a target with an initially unknown location as quickly as possible. To achieve this, control policies must exploit the diverse capabilities of the team, encouraging specialized behaviors for each agent class. For instance, explorer robots, being faster and more agile, are tasked with rapidly mapping the environment and locating the target. In contrast, rescuer robots, though slower, have the capability to engage with the target once its location is known.
To solve this problem, we propose a hierarchical control approach that combines high-level decision-making and low-level motion control for effective navigation. First, a goal location within a localized area centered on the robot is selected by the high-level decision-making module; then, the local planner generates motion commands to ensure smooth movement and obstacle avoidance en route to the goal.
@@
We formalize this approach as a Class-based Macro Action Decentralized Partially Observable Markov Decision Process (CMacDec-POMDP), a novel extension of the MacDec-POMDP framework [12] that incorporates varying properties across different agent classes. Assuming a team of $N$ agents with $C\leq N$ different classes, the problem is formalized as the tuple $\langle I,C,S,\{M^{c}\},\{A^{c}\},T,\{R^{c}\},\allowbreak\{\zeta^{c}\},\{Z^{c}\},\{\Omega^{c}\},\{O^{c}\},\gamma,h\rangle$, where
  • 1.
@@
  • 4.
    $M^{c}$ is a finite set of macro-actions (MAs) for agents of class $c$. The set of joint MAs is then $M=\times_{i}M^{C(i)}$;
  • 5.
@@
  • 8.
    $\zeta^{c}$ is a finite set of macro-observations (MOs) for agents of class $c$. The set of MOs is then $\zeta=\times_{i}\zeta^{C(i)}$;
  • 9.
@@
The solution to the CMacDec-POMDP is the joint high-level policy $\bar{\mu}=(\mu^{1},\dots,\mu^{N})$ that maximizes the expected cumulative discounted return from all agents,
@@
It should be noted that a generic parameter $p$ is denoted as $p^{i}$ when referring to the $i$-th agent, as $p^{i_{j:k}}=(p^{i_{j}},p^{i_{j+1}},\dots,p^{i_{k}})$ when referring to a subset $i_{j:k}$ of agents, and as $\bar{p}=(p^{1},p^{2},\dots,p^{N})$ when referring to the team of $N$ agents.
Figure 1: Workflow of the CATMiP framework during the execution phase. The robots communicate to share mission information and their embedded macro-observations to make high-level navigation decisions in a distributed manner, while a local motion controller generates the immediate action for each robot.
4 System Architecture
The overall architecture of the proposed Cooperative and Asynchronous Transformer-based Mission Planning (CATMiP) framework is shown in Figure 1. CATMiP is composed of three key modules: the C-SLAM module, which provides global mapping and localization; the High-Level Decision Making module, which determines macro-actions for agents and handles inter-agent communication; and the Low-Level Motion Control module, which translates macro-actions into motion commands for navigation. To enable efficient and scalable operation, CATMiP employs an asynchronous centralized training process to learn agent policies and an asynchronous distributed execution scheme to deploy these policies during real-time missions. The architecture is designed to leverage the diverse capabilities of heterogeneous agents while maintaining scalability to larger team sizes and environments.
4.1 Collaborative Simultaneous Localization and Mapping (C-SLAM)
This module enables robots to collaboratively build a shared global occupancy grid map. A distributed C-SLAM method that works under intermittent communication conditions, such as Swarm-SLAM [13], can be utilized during the execution phase. A designated robot temporarily processes data received from connected robots to generate the most accurate map and pose estimates, ensuring convergence to a unified global reference frame after a few rendezvous among subsets of robots. The occupancy grid map is then passed unto the high-level decision-making and low-level motion control modules.
4.2 High-Level Decision-Making
The High-Level Decision Making module determines macro-actions for each agent and enables coordination by leveraging information from robot sensory data, the C-SLAM module, and inter-agent communication.
During the centralized training phase, the Mission Coordinator sub-module aggregates mission-related information and forms all agents’ MOs, denoted as $\bar{z}$. The Mission Coordinator maintains the action-observation history $\bar{H}_{A}$ and the macro-action-macro-observation history $\bar{H}_{M}$. It manages agent activation and determines new MAs by passing $\bar{z}$ through the Asynchronous Multi-Agent Transformer (AMAT) network. The selected MAs are then sent to the Low-Level Motion Control module as navigation goals for execution.
@@ 4.2 High-Level Decision-Making
At each time step of the mission, whether during training or execution, agents are categorized as active, inactive, or on-standby, based on their current state and macro-action progress. Inactive agents continue executing primitive actions based on their current MA, as dictated by their low-level policy. Agents enter standby mode when their current MA termination condition is met. During training, agents on standby delay selecting their next MA for a random minimum number of time steps. This delay serves two purposes: it simulates the variable time required to complete MAs in real-world deployments and allows multiple standby agents to simultaneously activate and update their MAs with a single pass through the AMAT network. If an agent’s maximum standby time is reached and no other agents are on-standby, it activates alone. Agents that receive new MAs return to the inactive state to begin executing the newly assigned goal.
In the distributed execution phase, the Mission Coordinator operates locally on each robot, managing agent activation independently. It also facilitates the exchange of MOs and MAs between communicating agents. This localized approach ensures efficient real-time decision-making while maintaining coordination across the team.
4.3 Low-Level Motion Control
The Low-Level Motion Control module generates the robot’s primitive action $a^{i}_{t}$ at each time step $t$ by integrating a path planner and a motion controller. The path planner determines a collision-free path to the global goal specified by the robot’s current MA $m^{i}$. A path-finding algorithm, such as A* search [29], can be used to compute the shortest path on the occupancy grid map, represented as a sequence of waypoints.
To refine the planned path, the motion controller employs a local planner [30], which optimizes the robot’s trajectory based on the selected path and real-time sensory data. The local planner ensures that the resulting trajectory adheres to the robot’s motion constraints, avoids dynamic and static obstacles, and minimizes execution time. The optimized trajectory is then translated into low-level motion control commands, producing the primitive action $a^{i}_{t}$.
4.4 Asynchronous Centralized Training
@@
At every time step $t$ of a training episode, the agents execute the joint primitive action $\bar{a}_{t}$ generated by their low-level motion control module, and each collect a class specific state-action-dependent reward, jointly denoted as $\bar{R}(s_{t},\bar{a}_{t})$. At the $\tau$th instance of there being a subset $i_{1:g(\tau)}$ of agents ready to update their MAs (flagged as active agents), an ordered sequence of MOs $\bar{z}_{\tau}=\{z^{i_{1}}_{\tau},\dots,z^{i_{g(\tau)}}_{\tau},z^{i_{g(\tau)+1}}_{\tau},\dots,z^{i_{N}}_{\tau}\}$ are formed and fed as the input to the AMAT network. Subsequently, a sequence of latent representation of MOs of active agents, denoted as $\hat{z}^{i_{1:g(\tau)}}_{\tau}$, are calculated and used to obtain new MAs $m^{i_{1:g(\tau)}}_{\tau}$ and value function estimates $V(\hat{z}^{i_{1:g(\tau)}}_{\tau})$. The rewards accumulated by the active agents since their last activation, $R^{i_{1:g(\tau)}}_{acc}$, along with $\bar{z}_{\tau}$, $m^{i_{1:g(\tau)}}_{\tau}$, and $V(\hat{z}^{i_{1:g(\tau)}}_{\tau})$ are stored in a replay buffer. At the end of each training episode, the stored transitions are used to optimize the AMAT network parameters through backpropagation. The pseudocode of the centralized training process is presented in Algorithm 1.
The resulting fully trained network is then deployed for decentralized or distributed execution on each robot’s local hardware, enabling real-time inference of macro-actions during the mission.
@@
4.5 Asynchronous Distributed Execution
During the execution phase, the robots operate in a fully distributed manner, utilizing local instances of the AMAT network and time-varying communication neighborhoods formed through a dynamic mobile ad hoc network [31]. The communication topology is modeled as a dynamic graph $\mathcal{G}=(\mathcal{I},\mathcal{E})$, where vertices $\mathcal{I}$ represent the set of agents, and edges $\mathcal{E}$ represent communication links between them. The probability of a communication link $E_{ij}$ existing between agents $i$ and $j$ depends on their distance $d_{ij}$, defined as:
$Pr(E_{ij})=e^{-d_{ij}^{2}/\sigma^{2}},$ (2)
where $\sigma$ is a decay parameter controlling how quickly the communication probability decreases with distance. This probabilistic model is adapted from [32]. Each agent $i$ maintains a local communication neighborhood $\mathcal{N}(i)$, consisting of all agents (including itself) that can exchange information either directly or indirectly at a given time step.
Agents within the same communication neighborhood that become active at the same time step coordinate to elect a temporary broker. The broker facilitates local decision-making by aggregating neighborhood information and generating macro-actions for all active agents in its vicinity. It forms an ordered sequence of MO embeddings $(z_{e}^{i_{1}},\dots,z_{e}^{i_{g}},z_{e}^{i_{g+1}},\dots,z_{e}^{i_{m}})$, where $z_{e}^{i_{1}}$ corresponds to the broker’s own MO embedding, $z_{e}^{i_{2:g}}$ are embeddings received from other active agents in the neighborhood, and $z_{e}^{i_{{g+1}:m}}$ are the current MO embeddings from inactive agents in the neighborhood. The broker feeds this ordered sequence into its local instance of the AMAT network. The network outputs a sequence of updated MAs $(m^{i_{1}},\dots,m^{i_{g}})$ for all active agents in the neighborhood. The broker transmits the newly generated macro-actions back to the respective agents. These macro-actions serve as navigation goals and are passed to each robot’s local Low-Level Motion Control module to execute corresponding primitive actions.
5 Asynchronous Multi-Agent Transformer (AMAT)
This section details the structure and different components of the AMAT network and its use during asynchronous centralized training and distributed execution.
The AMAT network, illustrated in Fig. 2, consists of four components: Macro-Observations Embedder, Encoder, Macro-Actions Embedder, and Decoder. AMAT transforms a sequence of MOs $z^{i_{1:m}}$ from a subset $i_{1:m}$ of agents into a sequence of MAs $m^{i_{1:g}}$ corresponding to the active subset $i_{1:g}$ ($g\leq m$).
During centralized training, $i_{1:m}$ represents the complete agent set $I=\{1,\dots,N\}$, while in distributed execution, it refers to the agents in the neighborhood $\mathcal{N}(i)=\{i_{1},\dots,i_{m}\}$. In both cases, the sets are ordered to put the active agents $i_{1:g}$ first.